LHC 2011 - latest news

Week 14: Machine coordinators: Jan Uythoven and Jorg Wenninger

LHC coordination vistar




Link to report from morning meeting

Monday 11th April

  • 07:00: beam lost afte 10 Hz RF trim.

8:30 meeting 

Sunday 10th April

  • Dumped with 732/660 due to an RF arc (so said the alarm)
  • Achieved 804/804! but B2 was very unstable. Eventually dumped by the SIS
  • Stopped at 804/732. Stayed like that for ~45 minutes but the dumped by BPM in IP6
  • Refill for scrubbing with 800 bunches.
  • 12:15: 804b+804b. 1E14 protons in each beam.
  • 15:02: dump by RF interlock, arc.
  • 16:00: refill to 804b+804b.
  • 20:45: dump.
  • Loss maps.
  • 22:00: Green light to increase bunch limit to 1000 bunches.
  • 02:00: 1020 bunches circulating in each beam.

8:30 meeting 

Saturday 9th April

  • 10:30 Filled to 588 b / beam with 72b per injection and larger separation.
  • 14:30: beam dumped.
  • Loss maps. First loss maps shows again large hierarchy breaking in B1H. A second loss map after recalibration for all the BPMs is good.
  • 19:00: Green light to increase bunch limit to 800 bunches.
  • 20:00 :Filling.
  • 20:45: PC trip RQ5.L5 at 300b+300b.
  • Refill for 800 bunches

8:30 meeting

Friday 8th April

  • Verification of cabling between current lead voltage tabs and QPS by running special power cycles and checking the QPS logging signals.
  • Multiple accesses for QPS and EPC to restore powering.
  • All IPQs, IPDs, ITs, RBs and RQs have been checked.
  • 19:00: Green light for powering from MPE.
  • 22:00: finally pre-cycling...
  • 23:30: Beams back in the ring.
  • Problems with satellites in the PS and with XPOC in the LHC
  • 04:00: Ready to inject 36b per train, but dumps due to satellites at the end of the trains.
  • 06:00 New filling with 72b per injection and larger batch separation.
8:30 meeting

Thursday 7th April

  • Access for temperature sensor on DFB in S45, ventilation door, RF, experiments
  • 16:30: Decision to stop ALL powering. The cryo temperature sensor problem this morning led to a quench of the HTS current lead, because an erroneous temperature reading induced a closing of the valves. The analysis of the QPS data revealed that one of the QPS quench detection boards was not connected to the correct voltage tabs, thus removing the redundancy of the protection. The acccess revealed a cabling error between a patch panel and the QPS electronics that was not detected. No powering of the high current circuits until all cabling has been verified. The extend of the verifications must still be defined.
  •  Machine is idle over night.

8:30 meeting


Wednesday 6th April

  • 08:00-16:00: Injection setup for 50 ns, increasing # bunches per injection, max 400 bunches.
  • Scrubbing with up to 408+300bunches (limit to 600 b).
  • 23:30: Access to S81 for RB. Issue with ventilation door.
  • Scrubbing with up to 408+336 bunches (limit to 600 b).
  • 06:30: cryo problem on S45. DFB temperature sensor.

8:30 meeting


Tuesday 5th April

  • 08:00-16:00: Injection setup for 50 ns, up to 36 b / injection.
  • Systematic loss map studies with increased coupling and orbit changes.
  • Overnight scrubbing with 50 ns, up to 300 bunches.

8:30 meeting

Monday 4th April

  • 08:00-16:00: Injection of up to 96 bunches (3x24) with 75 ns beams. Injection and abort gap cleaning. Strong loss reduction thanks to 'deeper' scraping in the SPS.
  • 17:00-21:00 : TCTH.4L2.B1 alignment in IR2 end of squeeze and collisions. Loss map in collisions (H, V, dp/p < 0).
  • Trip of RB.78. Access.
  • Loss map at injection.
  • 06:00: start of scrubbing with 75 ns beam.

8:30 meeting